OpenStudy (anonymous):

Find the equation of the line tangent to the curve y = (x+3)/√x at its point of inflection. Help please. :)

OpenStudy (anonymous):

y=x^1/2+3x^(-1/2) y'=1/2 x^-1/2-3/2 x^(-3/2) y' = 0 at max/ min/ inflection

OpenStudy (anonymous):

from there its easy

OpenStudy (anonymous):

Thank you. :)